OpenStudy (anonymous):

Check my answer for medal and fan please Find the volume of a cone with a base radius of 5yd and a height of 6yd . Write the exact volume in terms of π , and be sure to include the correct unit in your answer.

OpenStudy (anonymous):

I got 157

OpenStudy (anonymous):

\[Volume =\frac{ 1 }{ 3 }\pi r^2h=\frac{ 1 }{ 3 }\pi* 5^2*6 ~cubic~yards=?\]

OpenStudy (anonymous):

I got 157 again

OpenStudy (anonymous):

so 157 yd^2

jimthompson5910 (jim_thompson5910):

what is 5^2 equal to?

jimthompson5910 (jim_thompson5910):

hints: 7^2 = 7*7 8^2 = 8*8 9^2 = 9*9 10^2 = 10*10 etc etc In general x^2 = x*x

OpenStudy (anonymous):

25

jimthompson5910 (jim_thompson5910):

now what is 25 times 6 equal to

OpenStudy (anonymous):

150

jimthompson5910 (jim_thompson5910):

now multiply that by 1/3

OpenStudy (anonymous):

50

jimthompson5910 (jim_thompson5910):

So you have computed \[\large \frac{1}{3}*5^2*6\] You found that \[\large \frac{1}{3}*5^2*6=50\]

jimthompson5910 (jim_thompson5910):

Now just place the 'pi' back in \[\Large \frac{1}{3}\pi*5^2*6=50\pi\]

jimthompson5910 (jim_thompson5910):

\(\Large 50\pi\) is the exact volume of this cone

OpenStudy (anonymous):

sooo 50 *3.14 = 157?

jimthompson5910 (jim_thompson5910):

when they say exact volume they mean in terms of pi. They even say "in terms of π" in the instructions. That means you do NOT replace pi with 3.14 and multiply. You just leave the pi alone. So it's simply \(\Large 50\pi\)

OpenStudy (anonymous):

OH!!!!

jimthompson5910 (jim_thompson5910):

157 would work for the approximate volume though

OpenStudy (anonymous):

thank you. so its 50pi yd^2?

jimthompson5910 (jim_thompson5910):

volume is always with "cubic" in it cubic feet cubic inches cubic millimeters etc

jimthompson5910 (jim_thompson5910):

yd^2 = square yards = area or surface area

OpenStudy (anonymous):

so its ^3

jimthompson5910 (jim_thompson5910):

yes yd^3 = cubic yards

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np